User name J. Singh
Log entry time 08:17:27 on August19,2003
Entry number 111879
keyword=owl shift summary
0000:0003 waiting for beam to come back & recover from septum quench
0003:0013 helium flow to our hall went away causing the magnets to trip off.
0013:0700 recovering from magnet offing. RQ1 was particularly difficult. Vince fixed it.
0800 NMR = 44.65% , EPR = 42.73%